OpenStudy (anonymous):

In the light of my 3 previous problems, this will be very easy to do. Let m,n be integers and p be an integer greater or equal to zero, then \[\left(m-\sqrt{n}\right)^p+\left(m+\sqrt{n}\right)^p \] is an integer.

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

All you need is this problem http://openstudy.com/users/eliassaab#/updates/54998a46e4b0b8a54fae714e

OpenStudy (anonymous):

yes the coefficient are alternative so the square roots disapears

OpenStudy (anonymous):

m-sqrt n is real let it be r m+sqrt n is real let be be s done

ganeshie8 (ganeshie8):

I think this should work when \(p\) is even : \[(m-\sqrt{n})^p + (m+\sqrt{n})^p = \sum \limits_{k=0}^{p/2} \binom{2p}{2k} m^{p-k}n^{k} \in \mathbb{N}\]

ganeshie8 (ganeshie8):

binomial thm is too irresistable to not think about ;p

OpenStudy (anonymous):

that remind me of number theory question (m-sqrt n )^p+ (m+ sqrt n)^p=0 mod something

OpenStudy (michele_laino):

Please, note that I think the solution can be given using the induction principle on integer p, namely: for p=1, we have: \[(m-\sqrt{n})+(m+\sqrt{n})= 2m\] so it is true, for p=2, we have: \[(m-\sqrt{n})^{2}+(m+\sqrt{n})^{2}=m ^{2}+n-2m \sqrt{n}+m ^{2}+n+2m \sqrt{n}=2(m ^{2}+n)\] so it is true, Now I suppose that your statement it is true for an integer p-1, so it is true also for integer p-2. namely: \[(m+\sqrt{n})^{p-1}+(m+\sqrt{n})^{p-1}\] is an integer. Now, for p, we note that: \[[(m-\sqrt{n})^{p-1}+(m+\sqrt{n})^{p-1}][(m+\sqrt{n})+(m-\sqrt{n})]=\] \[(m-\sqrt{n})^{p}+(m+\sqrt{n})^{p}+(m-\sqrt{n})^{p-1}(m+\sqrt{n})+\] \[+(m-\sqrt{n})(m+\sqrt{n})^{p-1}\] which is an integer by induction hipothesis. Now we can observe that:

OpenStudy (michele_laino):

\[(m-\sqrt{n})^{p-1}(m+\sqrt{n})+(m-\sqrt{n})(m+\sqrt{n})^{p-1}=\]

OpenStudy (michele_laino):

\[=(m-\sqrt{n})^{p-2}(m-\sqrt{n})(m+\sqrt{n})+(m+\sqrt{n})^{p-2}(m-\sqrt{n})(m+\sqrt{n})=\]

OpenStudy (michele_laino):

\[=[(m-\sqrt{n})^{p-2}+(m+\sqrt{n})^{p-2}](m-n)\] which is an integer by induction hypotesis, so our thesis is demonstrated!

OpenStudy (michele_laino):

ooops.. \[[...](m ^{2}-n)\]

OpenStudy (michele_laino):

our thesis is true because: \[(m-\sqrt{n})^{p}+(m+\sqrt{n})^{p}\] is expressed as difference between two integers!

OpenStudy (kainui):

Since m is an integer, it doesn't matter what power it's raised to. When -sqrt(n) is raised to an even power it removes the square root and becomes an integer and it doesn't matter. When -sqrt(n) is raised to an odd power it cancels out with its counterpart in the other term on the right and it doesn't matter. box proof thingy here.